### v2

# A Coordinated Routing Approach for Enhancing Bus Timeliness and Travel Efficiency in Mixed-Traffic Environment

Abstract
In this paper, we propose a coordinated routing strategy aimed at improving bus schedule adherence and enhancing travel efficiency for connected and automated vehicles (CAVs) operating within a mixed-traffic urban network. Our approach capitalizes on the existence of dedicated lanes for buses and CAVs, leveraging real-time traffic data to dynamically reroute CAVs in anticipation of congestion. By continuously monitoring traffic conditions on dedicated lanes and tracking the real-time positions of buses, we enable the system to proactively adjust CAV routes when potential interference with bus operations is detected. This coordination mitigates delays affecting transit services and reduces travel time for CAVs. We evaluate the proposed strategy through simulation studies conducted in the SUMO. The results demonstrate significant improvements in both transit reliability and CAV operational performance across a range of traffic conditions.
I Introduction
Connected and automated vehicles (CAVs) have emerged as a central element in the evolution of modern transportation systems. Advances in vehicle-to-vehicle, vehicle-to-infrastructure, and vehicle-to-everything communication technologies highlight the significant potential of CAVs in improving traffic efficiency, reducing energy consumption, and enhancing road safety [1, 2]. While the adoption of CAVs is projected to grow substantially over the coming decades, the transition towards a fully autonomous fleet is anticipated to unfold gradually [3, 4]. During this period, transportation networks will experience prolonged phases of mixed traffic, where CAVs and human-driven vehicles (HVs) coexist. Understanding how to enhance the traffic efficiency under increasing penetration rates of CAVs is therefore critical [5, 6, 7]. Extensive research and experimental studies have shown that heterogeneous traffic environments involving CAVs can significantly reduce travel times [8, 9, 10]. Among various approaches, implementing suitable right-of-way management is pivotal for ensuring the seamless integration of CAVs into existing transportation networks and for fully leveraging their benefits throughout this transition.
CAV dedicated lane (DL) has been proposed as an effective strategy to enhance the benefits that CAVs can bring to transportation systems [11]. By providing a separate driving environment, DLs enable CAVs to operate with shorter reaction times and smaller headways, thus increasing lane capacity by up to three times compared to general-purpose lanes (GPLs) [12]. In addition, DLs isolate CAVs from the mixed traffic network, offering a safer driving environment for both CAVs and HVs. This contributes to reducing vehicles’ lane changing and overtaking maneuvers, which are among the leading causes of collisions and delays in mixed traffic environments [13]. Despite the promising benefits of DLs, the high costs and limited feasibility of constructing new lanes in urban areas, due to substantial construction expenses and restricted urban space, often outweigh the advantages of CAV DLs [14]. In this context, converting an existing GPL into a CAV DL tends to be a more practical approach. However, it may lead to substantial delays for other types of vehicles using GPLs due to the reduced road capacity [15]. Recent research has shown that the overall system benefit from converting a GPL to a CAV DL may be lower than maintaining the original configuration [16].
To address the tradeoff inherent in implementing DLs for CAVs, a promising strategy has been proposed: the CAV and bus joint DL. Unlike conventional approaches that require constructing new lanes or displacing HVs, joint use of DLs by buses and CAVs repurposes existing DLs, which were originally designed for buses to ensure their on-time arrivals and maintain public transit efficiency [17]. While these lanes provide dedicated driving environments for transit operations, the reserved space cannot be taken by other non-transit vehicles, lowering the traffic efficiency [18]. Allowing CAVs to access these lanes can enhance lane utilization while maintaining transit reliability. In the existing literature, Luo et al. developed a framework for the optimal placement of CAV DLs to maximize benefits for all vehicles in the system [19]. Chen et al. explored a space-time modeling framework that demonstrates the mixed-use of DLs can enhance CAV travel efficiency and alleviate the overall traffic congestion [20]. A systematic framework investigating a variety of lane policies, such as exclusive automated vehicle lanes, regular vehicle lanes, and mixed-use lanes, was investigated in [12]. While the developed approaches provide new insights into the potential benefits of jointly using DLs, these methods remain limited in their ability to dynamically manage CAVs’ routing, particularly in coordinating vehicle behaviors across the network to balance traffic congestion reduction with maintaining reliable bus travel times, especially under fluctuating demand and a dynamic travel environment.
In this paper, we develop an efficient dynamic routing approach to coordinate CAVs traveling in a mixed traffic network, allowing for the DL sharing with buses. The goal is to prioritize the buses’ adherence to their scheduled timeliness while taking advantage of the joint DLs for CAVs to improve road capacity, reduce congestion, and enhance system-wide efficiency. By leveraging real-time traffic data, the proposed model enables estimation of potential tense flow on network edges and dynamically navigates the rerouting of CAVs to mitigate congestion on DLs. The main contributions of this work are as follows: (i) We model the rerouting problem of CAVs in a mixed traffic network where buses and CAVs jointly use DLs to fully take advantage of DLs while enhancing buses’ adherence to schedules. (ii) We incorporate a dynamic route condition estimation scheme in the rerouting framework, enabling real-time monitoring of traffic flow and timely adjustment of CAV routes when potential DL congestion is detected. Simulation performed in SUMO based on realistic traffic network data in San Francisco demonstrates the effectiveness of the proposed approach, indicating feasibility and significant improvements in traffic efficiency through coordinated use of joint DLs among buses and CAVs.
II Problem Statement
Figure 1: An example illustrating the road transportation network, where bus stations ($\mathcal{S}$) are shown by red nodes, intersections ($\mathcal{I}$) are shown in black, and the dedicated lanes are shown by red edges.
Consider an urban road network represented by a directed graph $G\!=\!(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes the set of nodes and $\mathcal{E}\!\subseteq\!\mathcal{V}\!\times\!\mathcal{V}$ denotes the set of directed edges. As illustrated in Fig. 1, the node set $\mathcal{V}$ is comprised of two subsets, denoted as $\mathcal{V}\!=\!\mathcal{S}\cup\mathcal{I}$, where $\mathcal{S}$ represents the set of bus stations and $\mathcal{I}$ represents the set of road intersections. Each directed edge $(v,v^{\prime})\in\mathcal{E}$ models a road segment from node $v$ to node $v^{\prime}$, with $v,v^{\prime}\in\mathcal{V}$. Consider each edge consisting of two lanes in the same direction, i.e., joint DL and GPL, where the GPL can be used by both CAVs and HVs, while DLs are reserved only for joint use by buses and CAVs.
Throughout this paper, we consider three types of vehicles traveling in the transportation network, including HVs, CAVs, and buses. The sets of HVs, CAVs, and buses within the monitoring time horizon are denoted by $\mathcal{N}^{\text{hv}}$, $\mathcal{N}^{\text{cav}}$, and $\mathcal{N}^{b}$, respectively. Accordingly, the set of all vehicles is represented as $\mathcal{N}\!=\!\mathcal{N}^{\text{hv}}\!\cup\!{\mathcal{N}^{\text{cav}}}\!\cup\!{\mathcal{N}^{b}}$. As illustrated in Fig. 2, buses use only DL and aim to arrive at each bus stop on time, following a specific timetable. In contrast, HVs can use only GPLs for travel purposes and route planning. With higher flexibility, CAVs can switch between joint DLs and GPLs to improve travel efficiency while minimizing interruptions to buses. Compared to employing only GPLs, driving on joint DLs enables CAVs to keep a smaller inter-vehicle distance, hence improving road capacity while avoiding disturbance caused by other HVs on GPLs.
Figure 2: Illustration of the general-purpose lane and dedicated lane.
For any bus $b\in\mathcal{N}^{\text{b}}$, let $P_{b}$ denote its fixed route, which is represented as a sequence of edges as follows:
$P_{b}=\big\{(v_{b,1},v_{b,2}),(v_{b,2},v_{b,3}),\dots,(v_{b,N_{b}-1},v_{b,N_{b}})\big\},$ (1)
where $v_{b,k}\!\in\!\mathcal{I}$ with $k\!=\!1,\dots,N_{b}$ denotes the $k$th road intersection along the bus route $P_{b}$, and $N_{b}$ represents the total number of intersections on the route. We refer to $v_{b,k}^{s}$ the $k$th bus stop on the edge $(v_{b,k},v_{b,k+1})$.
For each CAV $i\!\in\!\mathcal{N}^{\text{cav}}$ entering the road network, let $(o_{i},d_{i})$ denote its origin-destination (OD) pair, where $o_{i},d_{i}\!\in\!\mathcal{I}$. Based on real-time traffic conditions, CAV $i$ is assigned an optimal route from its location to the destination. The route of CAV $i$ from $o_{i}$ to $d_{i}$ is denoted by
$P_{o_{i},d_{i}}=\big\{(o_{i},v_{k}),(v_{k},v_{k+1}),\dots,(v_{n},d_{i})\big\},$ (2)
where $v_{k},v_{k+1},\dots,v_{n}\!\in\!\mathcal{I}$ are road intersections in the pre-planned route of the CAV. Note that, the route of each CAV is dynamically updated upon its arrival at each intersection in its route. While CAV routes may traverse both intersections and bus stations, re-optimization of the route toward the destination is triggered only when approaching an intersection. HVs are treated as passive participants in the network, following certain routes during their trips.
In this paper, we propose to develop a real-time coordinated routing approach for CAVs that fully leverages dedicated lanes, improves bus schedule adherence, and enhances CAV travel efficiency without creating new obstacles for timely bus arrivals at stops. To this end, a dynamic transportation system is modeled based on real-time traffic flow on each edge. Furthermore, CAVs are selected and rerouted to ensure on-time bus arrivals and maintain traffic efficiency for CAVs.
III Modeling of Traffic Flow
This section presents the modeling of road traffic flow in a dynamic transportation network. We start by introducing the traffic flow on joint DLs to capture the potential delays caused by CAVs. Then, the traffic flow on GPLs is modeled for effective rerouting of CAVs.
III-A Traffic Flow on Joint Dedicated Lane
The travel time for a vehicle to traverse an edge depends on the real-time traffic flow along the edge. As the number of vehicles traveling on an edge increases, traffic flow becomes heavier and travel times tend to rise due to congestion. This nonlinear relationship is well captured by the classic Bureau of Public Roads (BPR) function [21]. Let $\tau_{e}^{0}$ denote the free-flow travel time to traverse edge $e$. According to the BPR model, the actual travel time to traverse $e$ is denoted as
$\tau_{e}=\tau_{e}^{0}\left(1+\alpha\left(\frac{f_{e}}{c_{e}}\right)^{\beta}\right),$ (3)
where $f_{e}$ represents the real-time flow on edge $e$, and $c_{e}$ denotes the effective capacity of the same edge, i.e., the maximum flow that can be sustained without causing congestion. Both $f_{e}$ and $c_{e}$ are measured in the number of vehicles per unit time. In this paper, we assume that all vehicles, regardless of type, travel at the same speed under free-flow conditions. Additionally, $\alpha$ and $\beta$ are congestion sensitivity parameters. Since CAVs react faster and drive closer with shorter headways, they usually smooth traffic flow compared to a fully human-driven lane [19]. This results in a lower value of $\alpha$. In contrast, buses are heavier and less flexible, causing greater congestion when traffic is dense [22]. Based on these considerations, we adopt the following parameter settings in this study
$\begin{cases}\alpha\!=\!0.1,\ \beta\!=\!3,&\text{for GPL},\\ \alpha\!=\!0.2,\ \beta\!=\!5,&\text{for joint DL}.\end{cases}$ (4)
Sensors deployed in the road network enable real-time monitoring and analysis of traffic flow patterns [23]. In the road network considered, we assume that the traffic flow on each edge is measured by sensors located at the starting point of the edge. For each edge on the DLs of bus $b\!\in\!\mathcal{N}^{\text{b}}$, i.e., $(v_{b,k},v_{b,k+1})\!\in\!{P_{b}}$, the traffic flow monitor window is defined as
$\mathcal{M}(t_{\text{sys}},\Delta T)=\big[t_{\text{sys}}\!-\!\Delta T,\;t_{\text{sys}}\!+\!\Delta T\big],$ (5)
where $t_{\text{sys}}$ denotes the current system time and $\Delta{T}\in{\mathbb{N}}_{+}$ is a given constant representing the symmetric monitoring window where $\mathbb{N}_{+}$ denotes the set of positive integers. For each CAV $i\!\in\!\mathcal{N}^{\text{cav}}$ in the transportation system at time $t_{\text{sys}}$, let $P_{i}(t_{\text{sys}})$ denote its routing planning, which is represented as
$P_{i}(t_{\text{sys}})=\big\{(v_{i,1},v_{i,2}),(v_{i,2},v_{i,3}),\dots,(v_{i,n_{i}},d_{i})\big\},$
[…]
$t_{i,\ell+1}=t_{i,\ell}+\tau_{i,\ell}^{0}.$ (6)
This allows us to compute the accumulated number of CAVs expected to enter the edge $(v_{i,\ell+1},v_{i,\ell+2})$. To facilitate this, we define the following indicated function:
$I_{i}(t_{\text{sys}},\Delta T,v_{b,k})\!=\!\begin{cases}1,&\text{if }t_{i,\ell+1}\!\in\!\mathcal{M}(t_{\text{sys}},\Delta T),\\ &(v_{i,\ell+1},v_{i,\ell+2})\!=\!(v_{b,k},v_{b,k+1}),\\ \\ 0,&\text{otherwise}.\end{cases}$ (7)
Then, based on the number of CAVs expected to enter the edge $(v_{b,k},v_{b,k+1})$, the traffic flow on the $k$th edge in the DL of bus $b$ anticipated at time $t_{\text{sys}}$ is computed as
$\hat{f}_{b,k}(t_{\text{sys}})=\frac{\sum_{i\in\mathcal{N}^{\text{cav}}}I_{i}(t_{\text{sys}},\Delta T,v_{b,k})}{2\Delta T}.$ (8)
III-B Traffic Flow on General-Purpose Lane
Observing traffic flow on GPLs is essential to dynamically optimize CAV route planning based on real-time traffic data, thereby enhancing overall traffic efficiency. Similarly, the time window for traffic flow monitoring at the starting point of each edge on the GPLs is denoted as
$\mathcal{M}(t_{\text{sys}},\Delta\tilde{T})=\big[t_{\text{sys}}\!-\!\Delta\tilde{T},\;t_{\text{sys}}\!+\!\Delta\tilde{T}\big$], (9)
where $\Delta\tilde{T}\!\in\!\mathbb{N}_{+}$ represents the size of the time window to observe traffic flow. With the knowledge of the route planning of CAVs at time $t_{\text{sys}}$, the anticipated traffic flow on each edge belonging to GPLs is depicted by
$\hat{f}_{(v,v^{\prime})}(t_{\text{sys}})=\frac{\sum_{i\in\mathcal{N}^{\text{cav}}}\!I_{(v,v^{\prime})}(t_{\text{sys}},\Delta\tilde{T})+\Delta N^{\text{hv}}_{(v,v^{\prime})}(t_{\text{sys}})}{2\Delta{\tilde{T}}}$, (10)
[…]
$I_{(v,v^{\prime})}(t_{\text{sys}},\Delta\tilde{T})=\begin{cases}1,&\text{if }t_{i,\ell+1}\!\in\!\mathcal{M}(t_{\text{sys}},\Delta\tilde{T}),\\ &(v_{i,\ell+1},v_{i,\ell+2})\!=\!(v,v^{\prime}),\\ \\ 0,&\text{otherwise},\end{cases}$ (11)
where $(v,v^{\prime})\!\in\!\mathcal{E}\!\setminus\!\cup_{b\in\mathcal{N}^{\text{cav}}}P_{b}$ denotes an edge located on the GPLs. Note that $\Delta{N}^{\text{hv}}(t_{\text{sys}})$ represents the increase in the number of HVs, whose route choices are not controlled by the system planner but can be monitored.
Subsequently, according to the BPR model given in (3), the travel time for CAVs to traverse each edge on GPLs can be denoted by
$\hat{\tau}_{(v,v^{\prime})}(t_{\text{sys}})=\tau_{(v,v^{\prime})}^{0}\left(1+\alpha\left(\frac{\hat{f}_{(v,v^{\prime})}(t_{\text{sys}})}{c_{(v,v^{\prime})}}\right)^{\beta}\right).$ (12)
Here, $\tau_{(v,v^{\prime})}^{0}$ denotes the free-flow travel time on edge $(v,v^{\prime})$, and $\hat{f}_{(v,v^{\prime})}(t_{\text{sys}})$ is the anticipated traffic flow on the edge at time $t_{\text{sys}}$, computed according to (10). The congestion sensitivity parameters $\alpha$ and $\beta$ are determined by (4).
Remark 1
The traffic flow monitoring windows are defined by $\Delta{T}$ in (5) for joint DLs and by $\Delta{\tilde{T}}$ in (9) for GPLs, respectively, to capture real-time traffic conditions. A larger monitoring window puts greater emphasis on the macroscopic characteristics of traffic flow on network edges. With a different emphasis, the sizes of the two monitoring windows are not necessarily identical.
IV Dynamic Rerouting Approach
In this section, we propose a coordinated rerouting approach that dynamically optimizes route assignments for CAVs based on real-time traffic data. The objective is to ensure that buses adhere to their scheduled arrival times at each stop while enhancing the overall traffic efficiency of CAVs. The proposed optimization scheme consists of identifying the set of CAVs to be rerouted and then determining the optimal alternative routes for those CAVs.
IV-A Identification of CAVs for Rerouting
Let $\hat{t}_{b,k}$ denote the estimated arrival time of bus $b\!\in\!\mathcal{N}^{\text{b}}$ at its $k$th road intersection $v_{b,k}$. Upon arriving at its (k1)(k\!-\!1)th road intersection at time $t_{b,k-1}\!=\!t_{\text{sys}}$, the route ahead is denoted as
$P_{b}(t_{\text{sys}})=\big\{(v_{b,k},v_{b,k+1}),\dots,(v_{b,N_{b}-1},v_{b,N_{b}})\big\},$ (13)
excluding the intersection $v_{b,k-1}$. As such, the estimated arrival time $\hat{t}_{b,k}$ is represented as
$\hat{t}_{b,k}=t_{b,k-1}+\tau_{b,k-1}^{0},$ (14)
where $\tau_{b,k-1}^{0}$ denotes the free-flow travel time from intersection $v_{b,k-1}$ to intersection $v_{b,k}$.
Recall that the traffic flow $\hat{f}_{b,k}(t_{\text{sys}})$ on edge $(v_{b,k},v_{b,k+1})$ anticipated at time $t_{\text{sys}}$; cf. (8). Employing the BPR function in (3), the anticipated travel time on the same edge is as the following form:
$\hat{\tau}_{b,k}(t_{\text{sys}})=\tau_{b,k}^{0}\left(1+\alpha\left(\frac{\hat{f}_{b,k}(t_{\text{sys}})}{c_{b,k}}\right)^{\beta}\right),$ (15)
where $c_{b,k}$ is the effective capacity of edge $(v_{b,k},v_{b,k+1})$, defined as the maximum traffic flow that the road segment can accommodate under ideal conditions before congestion affects travel times. For edges on DLs, the parameters $\alpha$ and $\beta$ are chosen following the second case in (4).
Figure 3: Illustration of the rerouting scheme at each intersection in DLs.
To mitigate traffic congestion on each edge $(v_{b,k},v_{b,k+1})\!\in\!{P_{b}(t_{\text{sys}})}$ caused by CAVs, we define the control horizon for bus $b$ as the time interval $t_{b,k-1},\hat{t}_{b,k}]$. The traffic flow on edge $(v_{b,k},v_{b,k+1})$ is monitored starting from time $t_{b,k-1}$, and CAVs are dynamically rerouted to avoid their entry onto the edge $(v_{b,k},v_{b,k+1})$ whenever congestion is detected. We make use of Fig. 3 for better illustration.
For any system time $t_{\text{sys}}\!\in\![t_{b,k-1},\hat{t}_{b,k}]$, the set of CAVs on edge $(v_{b,k-1},v_{b,k})$ that need to be rerouted, denoted as $\mathcal{R}_{b,k-1}(t_{\text{sys}}^{\prime})$, is expressed as
$\mathcal{R}_{b,k-1}(t_{\text{sys}}^{\prime})=\big\{i\!\in\!{\mathcal{N}^{\text{cav}}}:I_{i}(t_{\text{sys}}^{\prime},\Delta{T},v_{b,k})\!=\!1\big\}.$ (16)
As defined (7), $I_{i}(t_{\text{sys}}^{\prime},\Delta{T},v_{b,k})\!=\!1$ indicates that CAV $i$ is within the monitor window of the intersection $v_{b,k}$ at time $t_{\text{sys}}^{\prime}$, which represents the time when congestion on edge $(v_{b,k},v_{b,k+1})$ is detected and CAV rerouting is triggered. Namely,
$t_{\text{sys}}^{\prime}\in\big\{t_{\text{sys}}\!\in\![t_{b,k-1},\hat{t}_{b,k}]:\hat{\tau}_{b,k}(t_{\text{sys}})\!\geq\!{\tau_{b,k}^{0}}\big\}.$ (17)
Remark 2
In the developed rerouting approach, the traffic flow on edge $(v_{b,k},v_{b,k+1})$ is dynamically monitored, and the CAVs on the preceding edge, i.e., $(v_{b,k-1},v_{b,k})$, are controlled before bus $b$ entering $(v_{b,k},v_{b,k+1})$. This procedure enables effective congestion mitigation on each edge along the DLs. Thus, it is reasonable to employ the free-flow travel time in (14) to estimate the bus’s arrival time at $v_{b,k}$.
IV-B Route Optimization for Rerouted CAVs
To mitigate traffic congestion on edge $(v_{b,k},v_{b,k+1})$ while improving the traffic efficiency of the transportation network, CAVs identified for rerouting are each assigned a new optimal route when $\hat{\tau}_{b,k}(t_{\text{sys}})\geq{\tau_{b,k}^{0}}$ is detected. Specifically, for each CAV $i\!\in\!\mathcal{R}_{b,k-1}(t_{\text{sys}}^{\prime})$, the system planner re-optimizes its route from intersection $v_{b,k}$ to the destination $d_{i}$, using the real-time traffic conditions observed at time $t_{\text{sys}}^{\prime}$. Based on the latest traffic data, including estimated travel times for each edge on GPLs and those on DLs, excluding the edge $(v_{b,k},v_{b,k+1})$, the optimization problem addressed for optimal route planning is cast as
$\min\ \sum_{i\in\mathcal{R}_{b,k-1}(t_{\text{sys}}^{\prime})}\sum_{(v_{i,\ell},v_{i,\ell+1})\in{P_{i}(t_{\text{sys}}^{\prime}})}\!\!\!\!\!\!\!\hat{\tau}_{i,\ell}(t_{\text{sys}}^{\prime})$ (18)
$\ \mathrm{s.\,t.}\quad~P_{i}(t_{\text{sys}}^{\prime})\!\in\!\mathcal{P}_{(v_{b,k},d_{i})},~i\in\mathcal{R}_{b,k-1}(t_{\text{sys}^{\prime}}),$ (19)
[…]
where $\hat{\tau}_{i,\ell}(t_{\text{sys}}^{\prime})$ represents the anticipated travel time to traverse edge $(v_{i,\ell},v_{i,\ell+1})$, estimated based on traffic conditions at time $t_{\text{sys}}^{\prime}$. As defined previously, $P_{i}(t_{\text{sys}}^{\prime})$ denotes the route planning of CAV $i$ at time $t_{\text{sys}}^{\prime}$ to complete the remainder of its trip. For CAVs to be rerouted, $P_{i}(t_{\text{sys}}^{\prime})$ refers to a route from intersection $v_{b,k}$ to the destination $d_{i}$. We denote by $\mathcal{P}_{(v_{b,k},d_{i})}$ the set of all feasible routes connecting $v_{b,k}$ to $d_{i}$.
Constraints (5)-(8), (14), and (15) provide estimates of traffic flow and travel times for each edge on joint DLs, while constraints (9)-(12) offer travel time estimates for edges on GPLs. Moreover, (16) and (17) specify the triggering condition for rerouting and identify the set of CAVs selected for route replanning. After all estimated travel times across the network are updated at time $t_{\text{sys}}^{\prime}$, a prediction-aware Dijkstra algorithm [24] is executed to determine the optimal rerouting paths for all identified CAVs, denoted as $P_{i}^{*}(t_{\text{sys}}^{\prime})$, where $i\!\in\!\mathcal{R}_{b,k-1}(t_{\text{sys}}^{\prime})$. In this modified Dijkstra framework, the edge weights correspond to the latest updated estimated travel times instead of static free-flow travel times or distances.
Remark 3
Solving the above route optimization problem requires real-time traffic data collected from existing sensor infrastructures, such as loop detectors and cameras. These data are aggregated within monitoring windows to compute vehicles’ actual travel times on the current edges and to estimate their travel times on future edges. Incorporating travel time estimations enables long-horizon congestion prediction and route adjustment in advance, thus preventing vehicles from being trapped in severe congestion.
Remark 4
The initial route of each CAV is computed based on the prevailing traffic conditions at the time the vehicle starts its trip, to minimize its own travel time. When rerouting is triggered for a group of CAVs, the system planner reoptimizes their routes to minimize their total travel time, thus enhancing overall traffic efficiency. This approach accounts for the diverse departure times of CAVs and ensures that the minimal set of CAVs is rerouted only when necessary, balancing CAVs’ flexibility, transit reliability, and system-wide efficiency.
V Simulation Study
This section presents simulation studies conducted on the Simulation of Urban MObility (SUMO) platform to evaluate the effectiveness and overall performance of the proposed approach in a mixed traffic network. We start by introducing the simulation setup and parameter configurations, and then present simulation results and analysis. The code implementation is available online. [^1] See code at: https://github.com/Tanlu-L/Coordinated-Routing-Approach-for-Enhancing-Bus-Timeliness-and-Travel-Efficiency
Figure 4: Illustration of the SUMO simulation environment, where the bus is depicted in yellow, CAVs are shown in red, and HVs are shown in blue. DLs are highlighted in grey, with bus stops marked in yellow.
TABLE I: Scheduled bus arrival times
Stop Station 1 Station 2 Station 3 Destination
Time [mm:ss] 01:03 03:46 07:39 09:27
V-A Simulation Setup
As shown in Fig. 4, we construct an urban road network in SUMO, which consists of $18$ road intersections and $3$ bus stops, with the road topology matching that shown in Fig. 1. This network replicates Van Ness Avenue in San Francisco from Market Street to Lombard Street, which includes one lane of DL. The roadway geometry and signal settings were obtained from OpenStreetMap. We consider that buses follow a predefined route along the joint DLs, traveling from the origin (node $7$ in Fig. 1) to the destination (node $15$). All other edges in the network are designated as GPLs. We assume sensors are installed at road intersections, enabling real-time observation of the number of vehicles entering each edge within a predefined monitoring time window. The monitoring time interval is set to $\Delta{T}\!=\!30$ seconds for joint DLs, and $\Delta\tilde{T}\!=\!60$ seconds for GPLs.
In the road network, we modeled the workday peak-hour bus service with $10$ buses operating along the same route from the origin stop to the destination stop, departing at intervals of $6$ minutes. The scheduled on-time arrivals at bus stops were determined based on free-flow travel times, with an additional $60$ seconds of dwell time at each station and a $30$-second disturbance margin to account for potential delays caused by traffic signals between stops. The scheduled arrival times for a representative bus are provided in Table I. CAVs and HVs enter the network at rates of $8$ and $20$ vehicles per minute, respectively, corresponding to a $30\%$ CAV penetration rate. The OD pair for each CAV is set as $(7,15)$, while the OD pair for each HV is randomly selected from the set $\{(1,6),(7,15),(16,21)\}$.
We compare the proposed coordinated routing strategy against two baseline methods: (i) static route planning (SRP) without rerouting and (ii) dynamic route planning (DRP) with rerouting. In the SRP strategy, routes are assigned based on free-flow travel times and remain fixed throughout the trip. In the DRP approach, each CAV is assigned the shortest path derived from real-time traffic data at the entry. Then, rerouting is applied if a lower-travel-time path is detected within a 60-second monitoring window. In the proposed method, CAVs follow the route recommendation obtained by the predictive control approach described in Section IV. Different from our approach, congestion anticipation on future road segments is not incorporated in the SRP and DRP. For both SRP and DRP, a modified Dijkstra algorithm using a weighted graph is applied to determine the shortest path. The weighted graph is initially constructed based on free-flow travel time on each edge, and the weights are updated dynamically according to the real-time traffic conditions.
Regarding the car-following-model in all three scenarios, CAVs use the intelligent driver model with a short headway ($0.6$ s) to capture automation-enabled efficiency. The maximum speed is set to $72\,\text{km/h}$, consistent with corridor speed limits. HVs and buses use the Krauss model: HVs with moderate headways (about $1.3$ s), variability to reflect human drivers, and a maximum speed of $72\,\text{km/h}$. Buses operate with larger headways ($1.6$ s) and a lower maximum speed of $65\,\text{km/h}$. This configuration reflects the capacity benefits of CAVs while realistically modeling human drivers and bus operations.
V-B Results and Analysis
Figure 5: The dynamics of the accumulated bus delay.
TABLE II: Bus on-time performance and delays at each station.
Station 1 Station 2 Station 3
On-time percentage [%]
   SRP using joint DL 10 30 30
   DRP using joint DL 50 60 60
   Proposed method 90 90 90
We evaluate the accumulated bus delay, defined as the total additional travel time experienced by all buses relative to their scheduled arrival times upon arrival at the destination, over a 3,600-second simulation horizon. The comparative results are presented in Fig. 5. As shown, buses experience significant delays under a static shortest-route assignment. The DRP approach reduces delays to some extent, but still, buses experience noticeable delays. In contrast, the proposed method significantly reduces bus delays nearly back to zero by integrating real-time traffic conditions and potential congestion estimation with efficient CAV rerouting. The results indicate that diverting CAVs away from congested edges, particularly those with bus stops during active dwell times, mitigates localized congestion along bus routes and improves schedule adherence. Table II reports the on-time percentage of $10$ buses at each station. The on-time percentage is defined as the fraction of buses arriving within $60$ seconds of their scheduled arrival times as given in Table I. It shows that the on-time percentages of the SRP, DRP, and our proposed method are approximately $23\%$, $57\%$, and $90\%$, respectively, demonstrating the superior performance of our method.
Figure 6: The dynamics of the cumulative CAV travel time.
Figure 7: The dynamics of the cumulative HV travel time.
Figure 8: The average bus delay at each stop.
Fig. 6 and Fig. 7 show the total travel time of CAVs and HVs over the simulation horizon. The proposed method achieves the lowest overall travel time for both CAVs and HVs. This improvement arises from proactively directing CAVs from congestion-prone edges, especially those affected by dwelling buses, thereby enhancing flow continuity throughout the network. Note that the DRP approach results in longer travel times for CAVs compared to both the SRP and the proposed method. Although for HVs, DRP initially reduces travel times, after $1800$ seconds, the travel times increase and even exceed those under SRP. These observations are common in transportation routing problems, as CAVs simultaneously identify and switch to shorter routes, which in turn leads to new and often more severe congestion on those routes.
Figure 9: The accumulated delay of each bus at experienced stops.
Fig. 8 shows the average trip delay for $10$ buses at each station, compared between the SRP method using the joint DL and the proposed approach. Bus stations indexed as $1$, 22, and $3$ correspond to nodes $9$, $11$, and $13$ in Fig. 1. As we can observe, the proposed method reduces bus delays significantly at all stations, achieving an average reduction of about $90.7\%$ per station compared with the SRP method. Fig. 9 further depicts the accumulated delay experienced by each bus across all stations along the route, showing that our rerouting strategy consistently reduces delays for individual buses at stop arrivals, thereby improving schedule adherence and overall transit reliability. Note that bus $6$ experiences relatively higher delays than others, mainly due to the great number of vehicles in the system. This result indicates that buses entering the network earlier or later experience less congestion. Such a pattern also aligns with the observed distribution of CAVs and HVs volumes in the system.
Figure 10: The average trip delay under different CAV penetration rates.
We further evaluate the efficiency of the proposed method under different CAV penetration rates. Fig. 10 presents the average trip delays experienced by buses, HVs, and CAVs under penetration levels of $10\%$, $30\%$, and $50\%$, with the total travel demand set to $2,000$ vehicles per hour. The results show that HV delays steadily decrease as CAV penetration increases, while CAV and bus delays exhibit little variation between $30\%$ and $50\%$ penetration. These findings show that the benefits of increased penetration saturate beyond a certain threshold, with limited additional travel time savings.
Figure 11: 90th-percentile travel time comparison with and without joint DL.
To further evaluate the efficiency of the proposed method, we introduce a scenario termed SRP without using joint DL. In this scenario, we replicate the setting of a real dedicated bus lane by allowing only buses to use the lane, thereby ensuring that buses experience no delay, while CAVs follow their initially assigned shortest routes. The results show that allowing CAVs to share the joint DL reduces travel times for both CAVs and HVs, but at the expense of significantly increasing bus travel times. In contrast, the proposed method maintains bus travel times close to those in the dedicated bus lane scenario, while further reducing CAV and HV travel times compared to SRP with joint DL.
Given the above results, we validate the effectiveness of the proposed rerouting strategy in preserving bus transit priority and improving schedule adherence. Through the joint use of DLs with buses, both CAVs and HVs benefit from the dynamic rerouting with reduced trip delays. These findings demonstrate that the congestion-aware coordinated routing approach improves travel efficiency across the entire system.
VI Conclusions
In this paper, we presented a coordinated routing strategy to improve bus schedule adherence and enhance travel efficiency for CAVs in a mixed-traffic urban environment. A central element of our approach was a real-time traffic monitoring system that estimates traffic flow conditions on each road segment using sensor data. Leveraging these estimates, we developed a coordination scheme that identifies a minimal subset of CAVs for rerouting and assigns them optimized paths reflecting not only current but also anticipated potential congestion. By dynamically updating CAV routes in response to time-varying travel environments along DLs, our method effectively mitigates disruptions to transit operations. Simulation experiments conducted in SUMO based on realistic urban road network data demonstrate that the proposed strategy improves both transit reliability and CAV performance under diverse traffic scenarios. Future work will focus on extending this framework to accommodate multi-modal transportation systems and assessing its scalability in large-scale urban networks.